OpenStudy (anonymous):

I GIVE GOLD MEDAL! given the domain values (-2, 0, 2) find the range for the function y=5x+1.

OpenStudy (anonymous):

Range are the y-values: So, when x = -2, y = 5(-2) + 1 = -9 When x = 0, y = 5(0)+1 = 1 When x = 2, y = 5(2) + 1 = 11 So the y-values, or range is {-9,1,11}.